# Crossover from Integer to Fractional Quantum Hall Effect

### v2

@@ Abstract
The parton theory constructs candidate fractional quantum Hall states by decomposing the physical particles into unphysical partons, placing the partons in integer quantum Hall states, and then gluing the partons back into the physical particles. Field theoretical formulations execute the gluing process through the device of emergent gauge fields. Here we study the process of going from the integer quantum Hall effect (IQHE) of fermionic partons to the fractional quantum Hall effect (FQHE) of bosons by introducing an attractive interaction between the fermions and continuously increasing its strength. It is far from obvious, a priori, that this process could be adiabatic, that is, go through without a gap closing, given that the wave function of the bosonic bound state and its Landau levels (LLs) involve all LLs of the fermions. The absence of an energy cutoff also makes the theoretical study challenging. We proceed by considering fermions on a lattice (a convenient lattice for our purposes is a subdivided icosahedron enclosing a magnetic monopole, which simulates the standard spherical geometry used in theoretical studies of the FQHE), wherein it is possible to work with the full Hilbert space during the entire process. For all finite systems that we have studied, we find that the system of two species of fermions occupying IQHE states evolves adiabatically into the FQHE of the bosons, suggesting that the process may be adiabatic also in the thermodynamic limit. This physics, reminiscent of the well studied BCS to BEC crossover, can in principle be realized in ultra-cold atomic systems consisting of two species of fermionic atoms in a synthetic magnetic field. Our study suggests that as the strength of the attractive interaction between the fermions of different species is increased, the system will exhibit a crossover, without gap closing, from two independent IQHE states of fermionic atoms to a FQHE state of bosonic molecules.
@@ I Introduction
The initial and the final states, i.e. $(1,1)$ and $1/2$, appear drastically different. First, while the initial $(1,1)$ state resides in the lowest Landau level (LLL) of the partons, the final 1/2 state involves all Landau levels (LLs) of the partons because of the delta function; i.e., the LLs of partons must strongly mix to form bosons and their LLs. Second, the initial and final states are topologically distinct. While the former has fermionic excitations, the latter has anyonic excitations. (We note that the charge of both excitations is one, but for fundamentally different reasons. For $(1,1)$ it is simply a charge one fermion, whereas for the $1/2$ state it is a fractionally charged excitation of charge-two bosons.) For these reasons, one might at first not suspect adiabatic continuity between the $(1,1)$ and $1/2$ states.
@@
We circumvent these problems by working on a lattice, which has a finite Hilbert space. We choose a subdivided icosahedron lattice with a radial magnetic field encoded in the hopping matrix elements that have Peierls phases associated with them. This simulates the standard spherical geometry that has been used in many numerical studies of the FQHE. This compact geometry without edges is favorable for demonstration of adiabatic continuity since it yields unique incompressible states with robust gaps. Using this setting, we numerically show that, at least for finite systems that we could study, the IQHE states of the partons are adiabatically connected, without any gap closing, to the bosonic FQHE states as we increase the strength of the attractive interaction.
While our model is motivated by the parton theory of the FQHE, it is interesting to ask if it can be realized by considering systems where partons are real (as opposed to fictitious) fermions. In principle, it applies to a system of spin-up and spin-down electrons each at an integral filling. In the limit when the electrons are non-interacting, both the up and down spin electrons fill an integer number of LLs, which is the case at even integer fillings in typical experiments. In the other limit with a strong onsite attraction, they exhibit a FQHE state of bosonic bound states consisting of two electrons. Achieving a strong attractive interaction between spin-up and spin-down electrons appears unrealistic, however. The same is true for a system of fermions in two layers. Perhaps a more promising platform is ultra-cold atomic systems. One can, in principle, achieve FQHE states in cold atom systems by creating a synthetic magnetic field through either rotation or complex hopping phases. We refer the reader to Ref. [32] for recent progress as well as for references to the earlier literature. One can thus imagine starting with the IQHE states of two species of noninteracting fermions (which could be the same fermionic atoms but in different internal states). By tuning through the Feshbach resonance it is possible to make the short range interaction between the different species of fermions dominant and attractive. Our study predicts that as the strength of the attractive interaction is enhanced, the system will evolve, without gap closing, into a FQHE state of bosonic molecules. This is a possible way to generate bosonic FQHE states in cold atom systems. The physics here is reminiscent of the BCS to BEC crossover [33, 34], which has been realized in ultra-cold Fermi gases [35, 36, 37]. Figure 1 illustrates the physics pictorially.
A notable feature is that while the ground state evolves adiabatically, the nature of the low-energy excitations changes qualitatively, from fermionic to anyonic, in the two limits. This may seem strange, but a change in the character of the excitations has been discussed for the transition between the Jain and the Gaffnian states at $\nu=2/5$ [38, 39, 40], as well as for the crossover between the BCS to the BEC state where the low-energy excitations go from broken pairs to density waves.
@@
For completeness, we also study this problem in the torus geometry, where a FQHE state is known to have degeneracies due to translation symmetry (and, additional degeneracies for non-Abelian states). In previous work studying an adiabatic process through anyons, the dimension of the Hilbert space changes discretely due to algebraic constraints of the braid group [41, 42], resulting in a peculiar structure in the adiabatic evolution: the ground state degeneracy changes wildly even though the energy gap varies smoothly [14]. This may be understood from the fact that apart from the degeneracy, the states have the same underlying topology. For our present case, we find that the IQHE state of partons evolves into one of the degenerate FQHE ground states of the physical particles, while other states approach it to produce the required degeneracy for the ground state in the $U=\infty$ limit. These additional states carry a nontrivial many-body Chern number, and the collection of degenerate states exhibits the same Hall conductance as the IQHE state of partons.
@@ III Hubbard Hamiltonian for partons
We investigate adiabatic continuity between the $(n_{\uparrow},n_{\downarrow})$ state and the ${n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}}$ state via exact diagonalization as we vary $U$ from 0 to a sufficiently large number. This article focuses on basic examples where $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,s;\frac{s}{1+s})$ with $s$ integer. The simplest case is $(1,1;1/2)$, which interpolates between the $(1,1)$ and $1/2$ states as $U$ is increased from 0 to $U/t\gg 1$. (Due to the contact interaction, the ground state of $H_{\text{eff}}$ at $\nu_{b}=1/2$ is the lattice analog of the Laughlin state.) Systems on an 1-icosahedron or on a torus are considered below. We will also consider the case of $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,2;2/3)$ and $(1,3;3/4)$ in the torus geometry. For exact diagonalization, we use the Lanczos method [48]. (We cannot access these systems on an 1-icosahedron. For the (1,2) state, say, we must have a minimum of $82N=$ 8 particles because it takes a minimum of 4 particles to fill the two LLs. The dimension of the Hilbert space for this system is 12,528,324,900 which is too large to perform exact diagonalization with computer resources currently available.)
@@ IV Numerical diagonalization
We begin with two species of fermions on a 1-icosahedron. Figure 5(a) shows the energy spectrum for $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,1,1/2)$ for a system of $42N=$ 4 fermions. (We measure all energies relative to the energy of the ground state.) As mentioned above, the $(1,1)$ ground state at $U=0$ is the tensor product of two IQHE states. The increase of $U$ from 0 leads to reconstruction of the single-particle spectrum from fermionic to bosonic LLs, which results in many level-crossings as shown in the figure. However, the ground state adiabatically evolves as $U$ is varied without any gap closing. While Fig. 5(a) shows the gap in the range of $U/t\in[0,20]$ (we quote energies in units of $t$), its inset plots the gap with $U/t\in[20,\infty)$ as a function of $t/U$. As $t/U$ decreases, the gap remains open in units of $\beta=2t^{2}/U$ and approaches the energy gap of the $\nu_{\text{b}}=1/2$ Laughlin state computed by diagonalizing the effective Hamiltonian $H_{\text{eff}}$ in Eq. (9). In Fig. 5(c), we show results for a system of $62N=$ 6 fermions. Due to large matrices, we plot only a few lowest energy states. Again, the energy gap of the $(1,1)$ state monotonically decreases and approaches the value estimated by the $\nu_{\text{b}}=1/2$ Laughlin state as $U$ increases. These results demonstrate adiabatic continuity between the $(1,1)$ state of partons and the bosonic $1/2$ Laughlin state for systems with 4 and 6 fermions. A study of larger systems is not possible due to the very large dimension of the Fock space (the dimension is 12,528,324,900 for $82N=$ 8 particles) but we think that is likely that the conclusion of adiabatic continuity remains valid also in the thermodynamic limit.
@@ IV.2 Torus
We consider a square lattice with $N_{x}\times N_{y}=6\times 6$ sites. Periodic boundary conditions are imposed in both directions unless otherwise stated. In Fig. 6(a), (b) and (c), the evolution of the lowest few eigenstates are shown for $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,1;1/2)$, $(1,2;2/3)$ and $(1,3;2/4)$. While the ground state is non-degenerate for small $U$, it has a degeneracy of 2, 3, and 4, respectively, at large $U$. The inset of Fig. 6(a) shows that, as $t/U$ decreases, the gap for the second-excited state scaled by $\beta=2t^{2}/U$ approaches to the value estimated by diagonalizing $H_{\text{eff}}$. The $(1,1)$ state “adiabatically” merges into the degenerate ground state sector at large $U$.
@@
For numerical calculation, we compute the discretized Berry curvature $\mathcal{F}(\vec{\eta})$ using the method in Ref. 50.
@@ V Concluding remarks
In this paper, we have developed an adiabatic scheme connecting IQHE and FQHE motivated by the parton construction of the FQHE. Specifically, we begin with an IQHE state of two fermionic partons and vary the attractive onsite interaction between partons from 0 to $\infty$ so they produce a FQHE state of bosonic bound states of partons. Because this involves a drastic reorganization of the states, it is necessary to include the full Hilbert space, and therefore we consider particles moving on a lattice. We consider both the spherical geometry, where the lattice is a subdivided icosahedron, and the torus geometry. Systems with 4 and 6 fermions are accessible to our study. In the spherical geometry, we find that the $(1,1)$ fermionic state adiabatically evolves into the 1/2 bosonic state. In the torus geometry, we find that the $(1,1)$, $(1,2)$ and $(1,3)$ fermionic states adiabatically merge into degenerate multiplets of the 1/2, 2/3 and 3/4 bosonic FQHE ground states. These small system studies suggest adiabatic continuity also in the thermodynamic limit. In contrast, there is no adiabatic continuity for the excited states. This observation — namely that the nature of excitations changes while the ground state evolves adiabatically – resembles the relation between the $\nu=2/5$ Jain state and the Gaffnian state [39, 40] or the BCS to BEC crossover. We have suggested that our model may be realized in ultra-cold atomic systems consisting of two species of fermions in a synthetic magnetic field, where the strength of the attractive interaction may be controlled by tuning through the Feshbach resonance.
@@ Acknowledgements.
We thank G. J. Sreejith and A. C. Balram for helpful discussions. K.K. acknowledges financial support from JSPS Overseas Research Fellowships. K.K. was supported in part by JSPS KAKENHI Grant no. JP23K19036 and JST CREST Grant no. JPMJCR18T2. J.S. was supported by the U. S. Department of Energy, Office of Basic Energy Sciences, under Grant no. DE-SC-0005042 and Award DE-SC0022245, and J.K.J. by the U.S. National Science Foundation under grant no. DMR-2037990. We acknowledge Advanced CyberInfrastructure computational resources provided by The Institute for CyberScience at The Pennsylvania State University and the computational resources offered by Research Institute for Information Technology, Kyushu University.
@@ Appendix A Effective Hamiltonian
where $\hat{n}_{i}=\hat{b}^{\dagger}_{i}\hat{b}_{i}$. The expression for a similar problem in Ref. [57] lacks the last term. By summing up $H_{\text{eff}}$ in Eq. (32) over all bonds for a given system, one obtains the general form shown in Eq. (9).

### v3

@@ Abstract
The parton theory constructs candidate fractional quantum Hall states by decomposing the physical particles into unphysical partons, placing the partons in integer quantum Hall states, and then gluing the partons back into the physical particles. Field theoretical formulations execute the gluing process through the device of emergent gauge fields. Here we study numerically the process of going from the integer quantum Hall effect of two species of fermionic partons to the fractional quantum Hall effect of bosons by introducing an attractive interaction between the fermions of different species and continuously increasing its strength to glue them into bosons. To properly capture the physics in the bulk, we implement this process in a lattice version of the spherical geometry, which allows us to keep the full Hilbert space. Even though the two end-point states are topologically distinct, we find that, for the small system sizes accessible to our study, the energy gap remains open, indicating a crossover between these two states.
@@ I Introduction
The initial and the final states, i.e. $(1,1)$ and $1/2$, appear drastically different. First, while the initial $(1,1)$ state resides in the lowest Landau level (LLL) of the partons, the final 1/2 state involves all Landau levels (LLs) of the partons because of the delta function; i.e., the LLs of partons must strongly mix to form bosons and their LLs. Second, the initial and final states are topologically distinct. While the former has fermionic excitations, the latter has anyonic excitations. [We note that the charge of both excitations is one, but for fundamentally different reasons. For $(1,1)$ it is simply a charge one fermion, whereas for the $1/2$ state it is a fractionally charged excitation of charge-two bosons.] For these reasons, one might expect the gap to close at some point as we go from the $(1,1)$ fermionic state to the $1/2$ bosonic state.
@@
We circumvent these problems by working on a lattice version of the spherical geometry, namely a subdivided icosahedron lattice, which has a finite Hilbert space. Many numerical studies of the FQHE have made use of the spherical geometry. We formulate a gauge convention to realize a uniform magnetic field on the subdivided icosahedron lattice. Using this setting, we numerically show that, at least for finite systems that we could study, the IQHE states of the partons are adiabatically connected, without any gap closing, to the bosonic FQHE states as we increase the strength of the attractive interaction.
The idea of going from a two-component integer quantum Hall state to a one component bosonic FQHE has been considered in the past [32, 33, 34, 35]. Repellin et al. have studied the problem on a lattice in the torus geometry and they find that the ground state becomes doubly degenerate at some value of the interaction, as expected given that the $1/2$ FQHE state is doubly degenerate on torus; they also identify a phase transition by considering the entanglement spectrum [34]. Yang and Zhai have derived the effective field theory of the quantum phase transition [32].
In the spherical geometry (a subdivided icosahedron lattice) used for our calculations, all incompressible FQHE states are non-degenerate. Also, the two terminal states, for example the $(1,1)$ and the $1/2$ states, occur at the same shift in the spherical geometry. The absence of a gap closing may seem inconsistent with the fact that the nature of the low-energy excitations changes qualitatively, from fermionic to anyonic, in the two limits. (Note that the fractional charge and statistics of the quasiparticle excitations can be defined for fairly small systems for the bosonic Laughlin 1/2 state [36, 37].) However, there exist other examples where the excitations change their character qualitatively even though the gap does not close (for small systems), such as the transition between the Jain and the Gaffnian states at $\nu=2/5$ [38, 39, 40] and the crossover between the BCS to the BEC state where the low-energy excitations go from broken pairs to density waves.
@@
For completeness, we also study this problem in the torus geometry, where a FQHE state is known to have degeneracies due to translation symmetry (and, additional degeneracies for non-Abelian states). Our results are consistent with those of Ref. 34. In previous work studying an adiabatic process through anyons, the dimension of the Hilbert space changes discretely due to algebraic constraints of the braid group [41, 42], resulting in a peculiar structure in the adiabatic evolution: the ground state degeneracy changes wildly even though the energy gap varies smoothly [14]. This may be understood from the fact that apart from the degeneracy, the states have the same underlying topology. For our present case, we find that the IQHE state of partons evolves into one of the degenerate FQHE ground states of the physical particles, while other states approach it to produce the required degeneracy for the ground state in the $U=\infty$ limit. These additional states carry a nontrivial many-body Chern number, and the collection of degenerate states exhibits the same Hall conductance as the IQHE state of partons.
While our model is motivated by the parton theory of the FQHE, it is interesting to ask if it can be realized by considering systems where partons are real (as opposed to fictitious) fermions. In principle, it applies to a system of spin-up and spin-down electrons each at an integral filling. In the limit when the electrons are non-interacting, both the up and down spin electrons fill an integer number of LLs, which is the case at even integer fillings in typical experiments. In the other limit with a strong onsite attraction, they exhibit a FQHE state of bosonic bound states consisting of two electrons. Achieving a strong attractive interaction between spin-up and spin-down electrons appears unrealistic, however. The same is true for a system of fermions in two layers. A more promising platform is ultra-cold atomic systems. One can, in principle, achieve FQHE states in cold atom systems by creating a synthetic magnetic field through either rotation or complex hopping phases. We refer the reader to Ref. 43 for recent progress as well as for references to the earlier literature. One can thus imagine starting with the IQHE states of two species of noninteracting fermions (which could be the same fermionic atoms but in different internal states). By tuning through the Feshbach resonance it is possible to make the short range interaction between the different species of fermions dominant and attractive. Our study suggests that as the strength of the attractive interaction is enhanced, the system with a finite number of particles will evolve, without gap closing, into a FQHE state of bosonic molecules. This is therefore a possible way to generate bosonic FQHE states in cold atom systems. The physics here is reminiscent of the BCS to BEC crossover [44, 45], which has been realized in ultra-cold Fermi gases [46, 47, 48]. Figure 1 illustrates the physics pictorially.
@@ III Hubbard Hamiltonian for partons
We investigate adiabatic continuity between the $(n_{\uparrow},n_{\downarrow})$ state and the ${n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}}$ state via exact diagonalization as we vary $U$ from 0 to a sufficiently large number. This article focuses on basic examples where $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,s;\frac{s}{1+s})$ with $s$ integer. The simplest case is $(1,1;1/2)$, which interpolates between the $(1,1)$ and $1/2$ states as $U$ is increased from 0 to $U/t\gg 1$. (Due to the contact interaction, the ground state of $H_{\text{eff}}$ at $\nu_{b}=1/2$ is the lattice analog of the Laughlin state.) Systems on an 1-icosahedron or on a torus are considered below. We will also consider the case of $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,2;2/3)$ and $(1,3;3/4)$ in the torus geometry. For exact diagonalization, we use the Lanczos method [54]. (We cannot access these systems on a 1-icosahedron. For the (1,2) state, say, we must have a minimum of $82N=$ 8 particles because it takes a minimum of 4 particles to fill the two LLs. The dimension of the Hilbert space for this system is 12,528,324,900 which is too large to perform exact diagonalization with computer resources currently available.)
@@ IV Numerical diagonalization
We begin with two species of fermions on a 1-icosahedron. Figure 5(a) shows the energy spectrum for $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,1,1/2)$ for a system of $42N=$ 4 fermions. (We measure all energies relative to the energy of the ground state.) As mentioned above, the $(1,1)$ ground state at $U=0$ is the tensor product of two IQHE states. The increase of $U$ from 0 leads to reconstruction of the single-particle spectrum from fermionic to bosonic LLs, which results in many level-crossings as shown in the figure. However, the ground state adiabatically evolves as $U$ is varied without any gap closing. While Fig. 5(a) shows the gap in the range of $U/t\in[0,20]$ (we quote energies in units of $t$), its inset plots the gap with $U/t\in[20,\infty)$ as a function of $t/U$. As $t/U$ decreases, the gap remains open in units of $\beta=2t^{2}/U$ and approaches the energy gap of the $\nu_{\text{b}}=1/2$ Laughlin state computed by diagonalizing the effective Hamiltonian $H_{\text{eff}}$ in Eq. (9). In Fig. 5(c), we show results for a system of $62N=$ 6 fermions. Due to large matrices, we plot only a few lowest energy states. Again, the energy gap of the $(1,1)$ state monotonically decreases and approaches the value estimated by the $\nu_{\text{b}}=1/2$ Laughlin state as $U$ increases. These results demonstrate adiabatic continuity between the $(1,1)$ state of partons and the bosonic $1/2$ Laughlin state for systems with 4 and 6 fermions. A study of larger systems is not possible due to the very large dimension of the Fock space (the dimension is 12,528,324,900 for $82N=$ 8 particles).
@@ IV.2 Torus
We consider a square lattice with $N_{x}\times N_{y}=6\times 6$ sites. Periodic boundary conditions are imposed in both directions unless otherwise stated. In Fig. 6(a), (b) and (c), the evolution of the lowest few eigenstates are shown for $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,1;1/2)$, $(1,2;2/3)$ and $(1,3;2/4)$. While the ground state is non-degenerate for small $U$, it has a degeneracy of 2, 3, and 4, respectively, at large $U$. The inset of Fig. 6(a) shows that, as $t/U$ decreases, the gap for the second-excited state scaled by $\beta=2t^{2}/U$ approaches to the value estimated by diagonalizing $H_{\text{eff}}$. These results are consistent with those of Ref. 34.
@@
For numerical calculation, we compute the discretized Berry curvature $\mathcal{F}(\vec{\eta})$ using the method in Ref. 56.
@@ V Concluding remarks
In this paper, we have developed an adiabatic scheme connecting IQHE and FQHE for systems with a finite number of particles, motivated by the parton construction of the FQHE. Specifically, we begin with an IQHE state of two species of fermionic partons and vary the attractive onsite interaction between partons from 0 to $\infty$ so they produce a FQHE state of bosonic bound states of partons. Because this involves a drastic reorganization of the states, it is necessary to include the full Hilbert space, and therefore we consider particles moving on a lattice. We consider both the spherical geometry, where the lattice is a subdivided icosahedron, and the torus geometry. Systems with 4 and 6 fermions are accessible to our study. In the spherical geometry, we find that the $(1,1)$ fermionic state adiabatically evolves into the 1/2 bosonic state. In the torus geometry, we find that the $(1,1)$, $(1,2)$ and $(1,3)$ fermionic states merge into degenerate multiplets of the 1/2, 2/3 and 3/4 bosonic FQHE ground states. In contrast, there is no adiabatic continuity for the excited states. This observation — namely that the nature of excitations changes while the ground state evolves adiabatically – resembles the relation between the $\nu=2/5$ Jain state and the Gaffnian state [39, 40] or the BCS to BEC crossover.
It is evident that the crossover from $(1,1)$ to $1/2$ state takes place at approximately $U\approx 10$. In Fig. 5, the energy gap separating the ground state and the first excited state changes its behavior qualitatively at $U\approx 10$. In the torus geometry, the ground state goes from a non-degenerate state to a nearly degenerate doublet at $U\approx 10$. Ref. 34 has identified a transition at this value by a consideration of the entanglement spectrum. Finally, Fig. 8 shows the ground state energy $E_{1}$, the first derivative $dE_{1}/dU$ (in the inset), and the second derivative $d^{2}E_{1}/dU^{2}$. While $E_{1}$ is a continuous function of $U$, $-d^{2}E_{1}/dU^{2}$ has a peak at $U\approx 10$, which becomes sharper as the particle number increases. Following the work of Wu et al. [35], this suggests a continuous phase transition in the thermodynamic limit.
Figure 8: (a) Ground state energy $E_{1}$ at $(n_{\uparrow},n_{\downarrow};{n_{\uparrow}n_{\downarrow}\over n_{\uparrow}+n_{\downarrow}})=(1,1;1/2)$ on a 1-icosahedron. The results in red and blue are for the systems considered in Figs. 5(a) and (b), respectively. The inset shows the first derivative $dE_{1}/dU$. (b) Second derivative $d^{2}E_{1}/dU^{2}$.
@@ Acknowledgements.
We thank Yinghai Wu, G. J. Sreejith and A. C. Balram for helpful discussions, and we are especially grateful to Yinghai Wu for bringing to our attention previous work along this direction. K.K. acknowledges financial support from JSPS Overseas Research Fellowships. K.K. was supported in part by JSPS KAKENHI Grant no. JP23K19036 and JST CREST Grant no. JPMJCR18T2. J.S. was supported by the U. S. Department of Energy, Office of Basic Energy Sciences, under Grant no. DE-SC-0005042 and Award DE-SC0022245, and J.K.J. by the U.S. National Science Foundation under grant no. DMR-2037990. We acknowledge Advanced CyberInfrastructure computational resources provided by The Institute for CyberScience at The Pennsylvania State University and the computational resources offered by Research Institute for Information Technology, Kyushu University.
@@ Appendix A Effective Hamiltonian
where $\hat{n}_{i}=\hat{b}^{\dagger}_{i}\hat{b}_{i}$. The expression for a similar problem in Ref. 63 lacks the last term. By summing up $H_{\text{eff}}$ in Eq. (32) over all bonds for a given system, one obtains the general form shown in Eq. (9).